factorize : 54x^3-16y^3
1
Expert's answer
2014-06-25T04:28:15-0400
Using formula a3-b3=(a-b)(a2+ab+b2) we obtain
54x3-16y3=2(27x3-8y3)=2(3x-2y)(9x2+6xy+4y2)